Three boxes of cookies are labelled A, B, and C. It is known that boxes A and B together contain 35 cookies, boxes A and C contain 28 cookies, and boxes B and C contain 37 cookies. How many cookies are in each box?
Answer by ankor@dixie-net.com(22740)

Three boxes of cookies are labeled A, B, and C.
It is known that boxes A and B together contain 35 cookies, boxes A and C contain 28 cookies, and boxes B and C contain 37 cookies.
How many cookies are in each box?
:
Add all three equations of each box
A + B + 0 = 35
A + 0 + C = 28
0 + B + C = 37
--------------
2A + 2B + 2C = 100
simplify divide by 2
A + B + C = 50
A + B + 0 = 35
----------------Subtract the 1st equation, eliminates A & B, find C
C = 15 in box C
:
A + C = 28
A + 15 = 28
A = 28 - 15
A = 13 in box A
:
B + C = 37
B + 15 = 37
B = 37 - 15
B = 22 in Box B
:
You can confirms these solutions in the given statements